OpenStudy (anonymous):

(6c+d/3c)-(7c+7d/4c)+2

OpenStudy (anonymous):

.Simplifying 5(c + d) + 2(3c + d) = 0 (c * 5 + d * 5) + 2(3c + d) = 0 (5c + 5d) + 2(3c + d) = 0 5c + 5d + (3c * 2 + d * 2) = 0 5c + 5d + (6c + 2d) = 0 Reorder the terms: 5c + 6c + 5d + 2d = 0 Combine like terms: 5c + 6c = 11c 11c + 5d + 2d = 0 Combine like terms: 5d + 2d = 7d 11c + 7d = 0 Solving 11c + 7d = 0 Solving for variable 'c'. Move all terms containing c to the left, all other terms to the right. Add '-7d' to each side of the equation. 11c + 7d + -7d = 0 + -7d Combine like terms: 7d + -7d = 0 11c + 0 = 0 + -7d 11c = 0 + -7d Remove the zero: 11c = -7d Divide each side by '11'. c = -0.6363636364d Simplifying c = -0.6363636364d

OpenStudy (anonymous):

C= -0.6363636364d

OpenStudy (anonymous):

thats not the answer im looking for